6x - 2y =1
10x -2y =5

Answer by mananth(16946)

6.00 x -2.00 y = 1.00 .............1
Total value
10.00 x -2.00 y = 5.00 .............2
Eliminate y
multiply (1)by -1.00
Multiply (2) by 1.00
-6.00 x 2.00 y = -1.00
10.00 x -2.00 y = 5.00
Add the two equations
4.00 x = 4.00
/ 4.00
x = 1.00
plug value of x in (1)
6.00 x -2.00 y = 1.00
6.00 -2.00 y = 1.00
-2.00 y = 1.00 -6.00
-2.00 y = -5.00
y = 2.50
x= 1.00
y= 2.50